a certain cnductor has a free electron density of 2x10^30 electrons per cubic feet. Its circular cross secctional area is 150000 square mils and 120 m long. Its resisitivity at 30 degrees. Celcius is 0.9x10^-8 ohm meter. If a current of 3 A passes thru this conductor. Find the drift velocty


1
Expert's answer
2021-10-07T23:42:17-0400

"V_d= \frac{I}{n *e*A}\\
V_d= \frac{3}{7629*10^{30} *1.6*10^{-19}*9.673*10^{-6}}\\
V_d=2.74*10^{-8} m/s"